3. An aqueous solution containing reactant A is fed to a batch reactor, operated as part of a continuous process involving other equipment. The feed rate to the reactor is 500 L/min. The reaction, A 2B C is second order irreversible, with the rate expression given as -ra 0.04CACB, ole/L min If C 1.7 gmole/L and C 3.0 gmole/L, what is the size of the reactor required to obtain an B0 80% conversion of A? The maintenance time between batches is 5 min.
Answer #1

Initial Conc of A = 1.7 gmol/L

Initial Conc of B = 3 gmol/L

Rate = 0.04CACB = 0.04 CA0 (1-x) (CB0 - CA0x) = 0.04 CA0 (1-x) CA0(M- x) = 0.1156 (1-x)(1.765-x)

Here M = CB0 / CA0 3/1.7 =1.765

For batch reactor

V0CA0 (dX/dt) = Rate x volume

V0CA0 integral ( 1/ rate ) dx = Volume x dt

(500 L/min x 1.7 gmol/L )* integral(1/ 0.1156 (1-x)(1.765-x) ) = Volume x integral(1 dt)

At t=0, x=0 and t=5 , x =0.8

Volume ,V = 437 L